# DriveRX: A Vision-Language Reasoning Model for Cross-Task Autonomous Driving

### v1

@@ 6 Conclusion
We propose AutoDriveRL, a reinforcement learning framework that decomposes autonomous driving into four reasoning subtasks, each guided by a task-specific reward model. This enables structured, interpretable reasoning and supports cross-task optimization.
The resulting model, DriveRX, achieves strong performance across perception, prediction, planning, and behavior tasks, outperforming GPT-4o on DriveBench behavior tasks and maintaining robustness under occlusion and other challenging conditions. Thanks to reward-guided reasoning, DriveRX produces shorter outputs and more efficient inference, better suited for real-time decision-making.
@@ Appendix A Limitations and Broader Impact
In this paper, we introduce AutoDriveRL, a unified training framework for vision-language models in autonomous driving that enhances reasoning and generalization across four core tasks; however, the approach still faces practical limitations.
Limitations. AutoDriveRL relies on manually designed, task-specific reward functions, which may limit its scalability to new tasks or sensor types. Additionally, the multi-stage inference pipeline introduces latency and has not been tested in real-time or closed-loop settings.
Broader Impact. By promoting interpretable, modular reasoning, AutoDriveRL may benefit the development of safer and more transparent autonomous systems. However, improper deployment of partially trained models or misaligned reward signals could lead to unsafe behavior. We encourage responsible use and further validation in realistic driving scenarios.
In the future, we will explore the effectiveness of AutoDriveRL across different frameworks and model sizes.
@@ Appendix C Data Composition
We conducted 8 inference runs for Align-DS-V, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-72B-Instruct on the DriveLM train dataset. Following the method in Eq. 1, we computed the difficulty of frames rather than individual questions. Figure. 4 shows the resulting difficulty distributions for each model. Next, we filtered out DriveLM samples with mutually exclusive options using specific rules and cleaned and balanced the dataset with methods like Action balancing.
To meet different requirements, we used Aligs-DS-V model scores to get frame scores, selected samples with filtered frame scores of $[25,45]$, and randomly sampled 1,000 frames as AutoDriveRL-Train. For the 3 models, we took samples with filtered frame scores of $[10,31]$, removed overlaps with the training set, and obtained DriveLM-Hard, which represents a more challenging subset of the data.
The task distribution of AutoDriveRL-Train and DriveLM-Hard can be observed in Figure. 5.
@@ Appendix D Models
LLaVA-1.5 [44] is an open-source multimodal vision-language model developed by the LLaVA team, designed to understand and generate language in response to visual inputs. It integrates a CLIP-based vision encoder with a Vicuna [56] language model through a multi-layer perceptron (MLP) connector, enabling it to perform tasks such as visual question answering, image captioning, and optical character recognition.
LLaVA-NeXT [45] is an advanced multimodal model based on LLaVA-1.5, developed to enhance performance in various applications. It demonstrates improved capabilities in reasoning, OCR, and world knowledge. The model supports higher image resolutions, utilizes refined datasets, and is built on more powerful language model foundations.
Qwen2.5-VL [47] is a multimodal model family developed by Alibaba Cloud, capable of handling tasks such as visual question answering, image/video captioning, and document understanding. It demonstrates strong performance across diverse vision-language applications.
InternVL3 [46] is a multimodal large language model family developed by Shanghai Artificial Intelligence Laboratory. It features a native multimodal pre-training approach, integrating text, image, and video modalities within a unified framework.
MM-Eureka [33] is a multimodal reasoning model extends large-scale rule-based reinforcement learning (RL) to multimodal reasoning.
R1-Onevision [37] is a multimodal reasoning model designed to bridge the gap between visual perception and deep reasoning.
Align-DS-V [41] is an experimental vision-language model that extends the DeepSeek-R1-Distill-Llama-8B, focusing on enhancing reasoning capabilities through all-modality alignment. It demonstrates strong performance on visual question answering and mathematical reasoning tasks.
DriveLM [35] is a vision-language model designed for end-to-end driving systems, developed by OpenDriveLab. It integrates large-scale vision-language models (VLMs) into driving systems to enhance generalization and interaction with human users.
Dolphin [36] is a multimodal vision-language model tailored for autonomous driving, built upon OpenFlamingo, and enhanced via Grounded Chain-of-Thought and in-context instruction tuning to achieve human-like understanding, reasoning.
GPT-4o [27] is a unified multimodal model by OpenAI that processes text, image, and audio inputs natively, enabling fast and accurate cross-modal reasoning with low latency.
@@ Appendix E Benchmark
DriveBench [34] is a benchmark dataset designed to assess the reliability of Vision-Language Models (VLMs) for autonomous driving. It consists of 19,200 frames and 20,498 question-answer pairs, covering a wide range of settings and tasks. It’s constructed by extracting a subset of frames from DriveLM-Train. Its innovation lies in being specifically created for evaluating the reliability of VLMs in autonomous driving scenarios, with diverse settings offering a comprehensive evaluation of VLMs’ performance in practical applications.
@@
DriveLM [35] is a method that uses a visual language model (VLM) to boost autonomous driving capabilities. It presents the Graph VQA task to replicate the multi-step reasoning of human drivers. Based on nuScenes and CARLA, the DriveLM dataset is created. The DriveLM-Agent, a VLM-based baseline approach, is proposed to perform Graph VQA and end-to-end driving jointly. Its innovation is introducing the Graph VQA task to integrate visual understanding, language interaction, and driving decisions, offering a new approach for autonomous driving development.
Appendix F Alignment Between Human and Model Rewards
@@
To reduce the cost of API-based evaluation during RL training and DriveLM-Hard benchmarking, we use Qwen2.5-72B-Instruct as the default reward model. To assess the validity of using open-source LLMs for reward estimation, we conduct a correlation analysis between model scores and human judgments.
@@
We design an online annotation platform for human evaluation and randomly select 300 samples from the evaluation set. Each sample is independently rated by three annotators. To ensure consistency, all annotators follow the same rubric used by our reward models, including task-specific criteria such as correctness, reasoning, and action relevance (see Appendix I). The average of the three scores is treated as the human gold score.
@@
Table 4: Pearson correlation between human scores and model scores.
@@
Appendix G Comparison with Different Training Strategies
To validate the effectiveness of joint rl in AutoDriveRL, we conduct comparative experiments under the same data scale. Specifically, we compare (i) supervised fine-tuning (SFT), (ii) task-wise separate training (on perception, prediction, planning, and behavior), and (iii) our joint RL training strategy. All three variants are based on the same base model, Align-DS-V, and are evaluated on DriveBench. The results are shown in Table 5
@@ Appendix G Comparison with Different Training Strategies
DriveRX-SFT $8$B $23.73$ $43.07$ $73.07$ $15.34$
DriveRX-Separate $8$B $24.73$ $46.59$ $76.11$ $55.18$
DriveRX-Joint $8$B $32.88$ $52.20$ $78.63$ $62.02$
In Table 5, we observe that the joint RL training scheme significantly improves performance across all four tasks under the same data scale. This result confirms that jointly optimizing multiple reasoning stages—rather than training them in isolation—enhances the model’s overall effectiveness in autonomous driving tasks.This highlights the benefit of jointly optimizing semantically distinct reasoning stages, allowing the model to learn shared representations and task interactions that improve generalization.
Appendix H Case Study
@@
Figure 11: Perception task example comparing DriveLM and DriveRX. The goal is to identify important objects in the current scene for future reasoning and driving decisions. DriveLM incorrectly includes irrelevant objects such as a fire hydrant and a bus (highlighted in red), while DriveRX correctly identifies the relevant vehicles in front of and behind the ego vehicle (highlighted in blue) by performing structured visual analysis. This perception task focuses on understanding key environmental elements that influence downstream planning and behavior.
@@
Figure 13: Planning task example comparing Align-DS-V and DriveRX. The goal is to determine safe actions for the ego vehicle in the current scene. Align-DS-V, while considering potential future obstructions, advises unnecessary caution due to the presence of a construction barrier and pedestrian, which may not be immediately relevant to the current driving situation (highlighted in red). In contrast, DriveRX provides a more accurate and practical recommendation by emphasizing maintaining the current speed on a clear road with no immediate hazards, as there are no visible pedestrians, other vehicles, or obstacles in the immediate vicinity (highlighted in blue). This planning task highlights the importance of analyzing visible obstacles and preparing for potential changes in road conditions to ensure safe driving, with DriveRX demonstrating a more reliable and contextually appropriate approach.
@@
Appendix I Experiment Prompt
I.1 Task-Specific Reward Prompt
@@ I.2 Inference Prompt
Appendix J Corruption Details

### v2

@@ 7 Conclusion
In this work, we proposed AutoDriveRL, a RL framework that decomposes autonomous driving into four interpretable reasoning subtasks—perception, prediction, planning, and behavior—each guided by a task-specific reward model. This design enables structured reasoning and improves decision interpretability. Based on this framework, we developed DriveRX, a reasoning-enhanced VLM that performs robust scene analysis across core driving tasks, achieving strong results and efficient inference under challenging conditions. Furthermore, we explored the downstream applications of reasoning-enhanced VLMs by implementing DriveRX-Agent for trajectory generation and DriveRX-VLA for action prediction. Both models achieve competitive performance in open- and closed-loop evaluations, demonstrating the feasibility and potential of a unified VLM backbone for supporting both high-level decision-making and low-level control.
8 Reproducibility statement
To ensure the reproducibility of our work, we will publicly release all relevant resources, including training and inference scripts, as well as the processed DriveLM-Hard dataset used in our experiments. The code will be made available via a link to a public repository. Additionally, we provide detailed descriptions of the data preprocessing steps, including the custom data filtering strategy applied to the DriveLM-Hard dataset, which is described in the supplementary materials. The experimental setup, including hardware, software versions, and model configurations, will also be documented in the appendix to facilitate accurate replication. We encourage other researchers to utilize these resources to reproduce and build upon our results.
@@ Appendix A Limitations and Broader Impact
In this paper, we introduce AutoDriveRL, a unified training framework for vision-language models in autonomous driving that enhances reasoning and generalization across four core driving tasks; however, the approach still faces practical limitations.
Reward Function Design. AutoDriveRL relies on manually designed, task-specific reward functions, which may limit its scalability to novel tasks or sensor modalities. Developing more generalizable or learnable reward signals remains an important direction for future research.
Efficiency Constraints. Inheriting the drawbacks of large language models (LLMs), our framework suffers from relatively long inference latency. Although our 8B-size model achieves performance comparable to or even surpassing 72B-scale models (see Table 1), the throughput remains a bottleneck, similar to prior works such as DriveLM (8.5 tokens/s in DriveLM-Agent). Unlike DriveLM, our approach does not require multi-round predictions, which leads to faster inference in practice. Moreover, recent advances in model optimization have improved LLM inference speed by orders of magnitude (Yang et al., 2025a; Bai et al., 2025), and we expect this trend to further alleviate the issue. We also experimented with a lightweight 3B VLA model, which significantly improves inference efficiency while maintaining competitive performance. In particular, the inference speed of Qwen2.5-VL-3B is approximately 2–3× faster than that of 8B models.
Applications and Scope. This section presents two application pathways of our framework — DriveRX-Agent for trajectory prediction and DriveRX-VLA for action-level control. These results show that injecting reasoning capabilities into a vision-language model (VLM) enables DriveRX to function as a high-level semantic reasoning backbone that provides structured cross-task understanding and intention reasoning for downstream planning and control models. Under the current experimental setup, both models achieve competitive performance in open- and closed-loop evaluations, though they still fall slightly short of the state-of-the-art (SOTA) on some metrics.
However, achieving the best low-level driving performance is not the primary goal of this work. Instead, our aim is to verify the effectiveness and applicability of reasoning-enhanced VLMs as a unified backbone for semantic perception, prediction, planning, and behavior abstraction in autonomous driving. While the structured reasoning outputs of DriveRX can benefit downstream planners and controllers, such extensions go beyond the scope of this paper and are left for future research. To this end, we adopt a simple and reproducible methodology, focusing on whether reasoning augmentation reliably improves high-level planning and semantic decision-making. Under this scope, DriveRX-VLA achieves competitive closed-loop results with limited data and minimal tuning, supporting our claim that a reasoning-based VLM can serve as a unified backbone for autonomous driving.
Reasoning Failure Cases. We observe that most failure cases arise in scenes where the Perception stage becomes unreliable due to severe occlusion or missing visual evidence. Since Perception initializes the entire reasoning chain, inaccuracies at this stage naturally propagate to subsequent Prediction, Planning, and Behavior stages, leading to inconsistencies across reasoning outputs. In addition, in a small number of low-confidence scenarios, DriveRX adopts a more conservative reasoning strategy, re-checking ambiguous cues and occasionally generating longer reasoning traces, which slightly increases the inference time. These behaviors occur only under highly uncertain visual conditions and do not affect overall performance or latency in normal cases.
Broader Impact. By promoting interpretable and modular reasoning, AutoDriveRL has the potential to advance the development of safer and more transparent autonomous systems. However, improper deployment of partially trained models or misaligned reward signals could lead to unsafe behavior. We encourage responsible use and further validation in realistic driving scenarios. Moreover, we plan to release DriveRX to the research community, hoping it can serve as a valuable source for distilling high-quality reasoning data and further accelerating progress in autonomous driving research. Importantly, the goal of this work is not to report improvements in real-time low-level driving performance, but to demonstrate the value of a reasoning-centered VLM backbone that unifies semantic perception, prediction, planning, and behavior abstraction. While the structured reasoning outputs of DriveRX can further benefit downstream planning or control modules, such extensions are beyond the scope of this work and are left for future research.
@@ Appendix C Data Composition
We conducted 8 inference runs for Align-DS-V, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-72B-Instruct on the DriveLM train dataset. Following the method in Eq. 1, we computed the difficulty of frames rather than individual questions. Figure 4 shows the resulting difficulty distributions for each model. Next, we filtered out DriveLM samples with mutually exclusive options using specific rules and cleaned and balanced the dataset with methods like Action balancing.
To meet different requirements, we used Aligs-DS-V model scores to get frame scores, selected samples with filtered frame scores of $[25,45]$, and randomly sampled 1,000 frames as AutoDriveRL-Train. For the 3 models, we took samples with filtered frame scores of $[10,31]$, removed overlaps with the training set, and obtained DriveLM-Hard, which represents a more challenging subset.
The task distribution of AutoDriveRL-Train and DriveLM-Hard can be observed in Figure 5.
@@ Appendix D Models
LLaVA-1.5 (Liu et al., 2024a) is an open-source multimodal vision-language model developed by the LLaVA team, designed to understand and generate language in response to visual inputs. It integrates a CLIP-based vision encoder with a Vicuna (Zheng et al., 2023) language model through a multi-layer perceptron (MLP) connector, enabling it to perform tasks such as visual question answering, image captioning, and optical character recognition.
LLaVA-NeXT (Liu et al., 2024b) is an advanced multimodal model based on LLaVA-1.5, developed to enhance performance in various applications. It demonstrates improved capabilities in reasoning, OCR, and world knowledge. The model supports higher image resolutions, utilizes refined datasets, and is built on more powerful language model foundations.
Qwen2.5-VL (Bai et al., 2025) is a multimodal model family developed by Alibaba Cloud, capable of handling tasks such as visual question answering, image/video captioning, and document understanding. It demonstrates strong performance across diverse vision-language applications.
InternVL3 (Zhu et al., 2025) is a multimodal large language model family developed by Shanghai Artificial Intelligence Laboratory. It features a native multimodal pre-training approach, integrating text, image, and video modalities within a unified framework.
MM-Eureka (Meng et al., 2025) is a multimodal reasoning model extends large-scale rule-based reinforcement learning (RL) to multimodal reasoning.
R1-Onevision (Yang et al., 2025b) is a multimodal reasoning model designed to bridge the gap between visual perception and deep reasoning.
Align-DS-V (PKU-Alignment, 2025) is an experimental vision-language model that extends the DeepSeek-R1-Distill-Llama-8B, focusing on enhancing reasoning capabilities through all-modality alignment. It demonstrates strong performance on visual question answering and mathematical reasoning tasks.
DriveLM (Sima et al., 2024) is a vision-language model designed for end-to-end driving systems, developed by OpenDriveLab. It integrates large-scale vision-language models (VLMs) into driving systems to enhance generalization and interaction with human users.
Dolphin (Ma et al., 2024) is a multimodal vision-language model tailored for autonomous driving, built upon OpenFlamingo, and enhanced via Grounded Chain-of-Thought and in-context instruction tuning to achieve human-like understanding, reasoning.
GPT-4o (OpenAI, 2024a) is a unified multimodal model by OpenAI that processes text, image, and audio inputs natively, enabling fast and accurate cross-modal reasoning with low latency.
@@ Appendix E Benchmark
DriveBench (Xie et al., 2025) is a benchmark dataset designed to assess the reliability of Vision-Language Models (VLMs) for autonomous driving. It consists of 19,200 frames and 20,498 question-answer pairs, covering a wide range of settings and tasks. It’s constructed by extracting a subset of frames from DriveLM-Train. Its innovation lies in being specifically created for evaluating the reliability of VLMs in autonomous driving scenarios, with diverse settings offering a comprehensive evaluation of VLMs’ performance in practical applications.
@@
DriveLM (Sima et al., 2024) is a method that uses a visual language model (VLM) to boost autonomous driving capabilities. It presents the Graph VQA task to replicate the multi-step reasoning of human drivers. Based on nuScenes and CARLA, the DriveLM dataset is created. The DriveLM-Agent, a VLM-based baseline approach, is proposed to perform Graph VQA and end-to-end driving jointly. Its innovation is introducing the Graph VQA task to integrate visual understanding, language interaction, and driving decisions, offering a new approach for autonomous driving development.
Bench2Drive (Jia et al., 2024) is a benchmark for evaluating end-to-end autonomous driving systems in closed-loop scenarios. It introduces 220 short routes covering 44 interactive driving scenarios, including varying weather and environmental conditions. Based on CARLA, the dataset provides 2 million annotated frames for fair and comprehensive evaluation. Bench2Drive allows for detailed testing of driving capabilities such as trajectory prediction and action execution, enabling the development of more robust autonomous driving models.
DriveAction (Hao et al., 2025) is the first action-driven benchmark specifically designed for Vision-Language-Action (VLA) models in autonomous driving. It comprises 16,185 QA pairs from 2,610 real-world driving scenarios. The benchmark uses driver-contributed real-world data and high-level discrete action labels collected directly from real-time driver operations. Its innovation is the action-rooted, tree-structured evaluation framework, which systematically links vision, language, and action tasks (V-L-A). This framework is used to rigorously evaluate state-of-the-art Vision-Language Models (VLMs) across different V-L-A modes, revealing that both vision and language guidance are necessary for accurate action prediction.
Appendix F Impact of Vision Encoders on Perception Tasks
Second, the visual encoder plays a critical role in determining the upper bound of perception task performance during training. In our experiments, the improvement of DriveRX over its base model Align-DS-V is limited (+1.47+1.47), despite a substantial gain on the behavior task (+11.49+11.49). As shown in the training curves (Figure 2(d)(b)), the perception reward for DriveRX saturates early (within 50 steps) and remains below 50, indicating a clear performance ceiling. We attribute this to the outdated architecture of CLIP-ViT-L/14, an early-generation ViT model whose limited perceptual capability falls behind modern ViT-based encoders.
However, high-performing models like Qwen2.5-VL (Team, 2024) and InternVL3 (Zhu et al., 2025) typically adopt dynamic resolution mechanisms, which allow them to process high-resolution inputs more effectively. While this improves visual understanding and perception scores, it also leads to significantly longer image token sequences. For instance, Qwen2.5-VL generates substantially more vision tokens per image compared to fixed-resolution VLMs like Align-DS-V. This increase in token length introduces notable inference latency, posing challenges for real-time decision-making in autonomous driving. In contrast to these models, DriveRX incorporates a fixed-length feature processing pipeline in its architectural design. While this design sacrifices some perceptual performance due to the output dimensionality constraints of the CLIP encoder, it achieves three critical advantages for practical deployment: deterministic inference efficiency, faster processing speed, and more predictable inference behavior.
Table 5: Impact of different vision encoders on the perception task. We report DriveBench perception and behavior scores (higher is better), average response length (in tokens) per question, number of image tokens per image, and inference time per question (in seconds). For a fair comparison, inference time is reported only among models of the same size. DriveLM requires 4 inference steps.
Model Size [Uncaptioned image] Vision Encoder [Uncaptioned image] Response [Uncaptioned image] Image Tokens [Uncaptioned image] Perception ($\uparrow$) [Uncaptioned image] Behavior [Uncaptioned image] Infer times
$M$-Eureka $7\text{B}$ Qwen-ViT-L/14 $543.12$ $1824$ $34.44$ $49.40$ $2.52$
R1-Onevision $7\text{B}$ Qwen-ViT-L/14 $447.55$ $1824$ $28.64$ $41.57$ $2.59$
Qwen2.5-VL $7\text{B}$ Qwen-ViT-L/14 $357.85$ $1824$ $31.92$ $47.40$ $2.76$
Qwen2.5-VL $72\text{B}$ Qwen-ViT-L/14 $343.27$ $1824$ $37.56$ $55.57$ -
InternVL3 $8\text{B}$ InternViT-300M-448px $270.94$ $2304$ $33.21$ $51.47$ $1.58$
InternVL3 $78\text{B}$ InternViT-6B-448px $275.69$ $2304$ $39.44$ $50.73$ -
DriveLM $7\text{B}$ CLIP ViT-L/14 $623.04$ $576$ $16.85$ $42.78$ $3.18$
Align-DS-V $8\text{B}$ CLIP ViT-L/14 $252.10$ $576$ $31.41$ $50.53$ $0.68$
DriveRX $8\text{B}$ CLIP ViT-L/14 $363.44$ $576$ $32.88$ $62.02$ $1.06$
Appendix G Ablation Study
G.1 Ablation of Subtasks
To further verify the effectiveness of joint training, we conducted additional ablation studies on the DriveLM-Hard benchmark by withholding training data for each individual subtask. As shown in Table 6, removing any of the four subtasks (Perception, Prediction, Planning, or Behavior) consistently degraded the final behavior score, highlighting that each task provides indispensable information for downstream decision-making. We also evaluated a sequential pipeline where the output of each task is directly cascaded into the next without joint training, and found that its performance was similarly unsatisfactory. This suggests that the benefit of our framework arises not only from the decomposition itself but also from the integrated optimization across subtasks, as the interdependence between earlier and later tasks is crucial for accurate reasoning.
G.2 Ablation of Data Filtering
Additionally, we include a controlled ablation to isolate the effect of our data-filtering strategy. Specifically, we randomly sample from the raw dataset the same number of training instances as in the filtered dataset, keeping all training configurations identical, and evaluate on DriveLM-Hard.
As shown in Table 7, the model trained on the filtered corpus outperforms the unfiltered baseline across all four core tasks on DriveLM-Hard. This consistent performance advantage across tasks demonstrates that our data-filtering strategy provides cleaner, more task-aligned supervision signals, which in turn enhance training stability and final model performance in RL.
Furthermore, the score curves on the validation set during training highlight the impact of data filtering. As shown in the figure 7 below, the model trained on the filtered data (blue curve) exhibits (i) much faster improvement on the validation set from the very first steps, and (ii) consistently higher performance compared to the model trained on randomly sampled data (red curve). The test curves, evaluated using the same test set, clearly illustrate the faster convergence and superior performance of the filtered-data model, emphasizing the advantages of our data-filtering strategy.
Figure 7: Score curves on the validation set during training.
Table 6: Ablation study on the DriveLM-Hard benchmark. We report task-wise VQA scores for each setting. Removing any subtask or replacing joint training with sequential cascading leads to clear performance degradation.
Model Perception Prediction Planning Behavior
w/o Perception 25.13 57.81 37.12 27.15
w/o Prediction 27.41 42.12 41.25 30.95
w/o Planning 27.21 60.43 34.05 31.20
Sequential Cascade 29.45 49.45 26.53 29.45
Align-DS-V 29.45 52.90 29.15 32.48
DriveRX (Ours) 34.83 71.84 51.42 36.82
Table 7: Ablation Study on the Effect of Data Filtering in RL Training. We compare the performance of DriveRX trained with filtered and unfiltered data across multiple tasks on the DriveLM-Hard. The table highlights the impact of data filtering on model performance and stability.
Model Perception Prediction Planning Behavior
Align-DS-V (base) 29.45 52.90 29.15 32.48
DriveRX w/ Data Filtering 34.83 71.84 51.42 36.82
DriveRX w/o Data Filtering 30.29 51.82 48.23 33.07
Appendix H Data Distillation Details
All training data for both DriveRX-Agent and DriveRX-VLA are distilled from the DriveRX reasoning corpus. We start from two original datasets processed by DriveLM (Sima et al., 2024): the nuScenes (Caesar et al., 2020) dataset for real-world scenes and the CARLA (Dosovitskiy et al., 2017) dataset for simulation. For each sample, we input the original prompt and ground-truth answer into DriveRX to generate an intermediate reasoning trace encapsulated in the field, which provides structured decision-making information beyond the final output.
Based on this reasoning-enriched data, we construct two types of supervision signals. For DriveRX-Agent, the ground-truth vehicle trajectory is projected into a discrete vocabulary space by partitioning continuous coordinates into 512 bins, each represented by a unique token. These trajectory tokens serve as supervision targets, enabling the model to directly output discrete waypoint sequences conditioned on camera observations, ego-vehicle states, and context. Case study in figure 17.
For DriveRX-VLA, we cluster short motion segments using a K-means algorithm to obtain 2048 discrete action tokens, each representing a short-term executable maneuver such as steering, acceleration, or braking. These tokens are incorporated into the model’s vocabulary, enabling the VLA model to directly predict control actions during inference.
During inference, both models receive multi-modal inputs, including camera images and ego-vehicle state information, and generate either trajectory tokens (DriveRX-Agent) or action tokens (DriveRX-VLA). These tokens are then decoded into continuous trajectories or executable control signals. This unified data distillation and tokenization pipeline ensures that both models learn from structured reasoning traces and ground-truth supervision, providing strong generalization and control capabilities.
Appendix I Alignment Between Human and Model Rewards
@@
To reduce the cost of API-based evaluation, we use Qwen2.5-72B-Instruct as the default reward model. To assess the validity of using open-source LLMs for reward estimation, we conduct a correlation analysis between model scores and human judgments.
@@
We design an online annotation platform for human evaluation and randomly select 300 samples from the evaluation set. Each sample is independently rated by three annotators. To ensure consistency, all annotators follow the same rubric used by our reward models, including task-specific criteria such as correctness, reasoning, and action relevance (see Appendix M). The average of the three scores is treated as the human gold score.
@@
The GPT-4o correlation (r = 0.83) demonstrates a strong alignment between model-based reward estimation and human judgment. Notably, this level of agreement matches the inter-annotator consistency reported for nuScenes-AP@0.5: when two human annotators draw bounding boxes, AP@0.5 considers them a match if their boxes overlap by at least 50% IoU (nuScenes (Caesar et al., 2020), App. A4). In other words, GPT’s judgment is about as consistent with a human grader as two human annotators are with each other under that widely accepted 0.5-IoU criterion.
Table 8: Pearson correlation between human scores and model scores.
@@
Appendix J More Evaluation Metrics
Figure 12: Training curves of GPT-based vs. hybrid reward scoring.
Evaluation of autonomous driving reasoning systems must be highly task-specific, as each sub-task emphasizes distinct aspects of perception and decision-making. In our framework, we design four dedicated evaluation prompts, corresponding to Perception, Prediction, Planning, and Behavior, ensuring that each score measures the attributes most relevant to its target task. For example, Perception requires not only identifying critical objects but also providing their spatial locations in structured coordinate form.
Conventional metrics often fail to capture the full complexity of such tasks. IoU-based metrics become unreliable when the predicted and reference object lists differ in number or ordering, and n-gram-based text similarity measures (e.g., BLEU, ROUGE) penalize valid paraphrases that nonetheless convey the correct semantics. These limitations motivate the use of GPT-based scoring, which can jointly consider semantic correctness, reasoning quality, and structural alignment in a unified evaluation.
To further investigate alternative evaluation strategies, we designed a hybrid reward scheme that integrates rule-based, semantic, and structural metrics, tailored to different question types:
  • Classification tasks: For questions with discrete options or yes/no answers, we apply rule-based scoring, assigning $1$ point for a correct match with the ground truth and 0 otherwise.
  • Open-ended tasks: For free-form answers, we compute semantic similarity between predictions and references using GPT-based scoring, capturing meaning beyond surface form.
  • Structured text + coordinate tasks: We separately evaluate textual descriptions with BLEU and ROUGEL metrics, while assessing coordinate precision independently.
  • Coordinate-only tasks: We extract all floating-point numbers from the predicted and reference outputs, group them into coordinate pairs, and compute an F1 score by considering a prediction correct if the absolute difference between corresponding coordinates is less than $16$ pixels.
Although the hybrid approach provides more fine-grained supervision, it introduces significant instability during reinforcement learning. Empirically, we observe that reward curves fluctuate heavily and are difficult to stabilize, in contrast to the smooth and monotonic improvements achieved with pure GPT-based scoring. Such instability may result from inconsistencies in metric scales, conflicts between optimization objectives, or the high sensitivity of structured matching signals to small output variations. As shown in Figure 12, even the best-performing hybrid runs exhibit large oscillations and lack a clear convergence trend.
These findings highlight a fundamental trade-off between metric granularity and training stability. Hybrid metrics offer more detailed signal decomposition but are highly susceptible to gaming behavior. GPT-based evaluation, by contrast, provides a holistic and robust measurement of task performance, balancing semantic understanding, structured reasoning, and action relevance. For this reason, we adopt GPT-based scores as our primary evaluation signal across all four task categories. Future work could explore dynamic weighting or staged evaluation to balance granularity and stability.
Appendix K Comparison with Different Training Strategies
To validate the effectiveness of joint rl in AutoDriveRL, we conduct comparative experiments under the same data scale. Specifically, we compare (i) supervised fine-tuning (SFT), (ii) task-wise separate RL training (on perception, prediction, planning, and behavior), and (iii) our joint RL training strategy. All three variants are based on the same base model, Align-DS-V, and are evaluated on DriveBench. The results are shown in Table 9
@@
DriveRX-SFT 88B $23.73$ $43.07$ $73.07$ $15.34$
DriveRX-Separate 88B $24.73$ $46.59$ $76.11$ $55.18$
DriveRX-JointRL 88B $32.88$ $52.20$ $78.63$ $62.02$
In Table 9, we observe that the joint RL training scheme significantly improves performance across all four tasks under the same data scale. This result confirms that jointly optimizing multiple reasoning stages—rather than training them in isolation—enhances the model’s overall effectiveness in autonomous driving tasks.This highlights the benefit of jointly optimizing semantically distinct reasoning stages, allowing the model to learn shared representations and task interactions that improve generalization. Furthermore, Table 9 shows that, on the behavior task, the performance of the model trained with SFT is significantly lower than that of the model trained with RL. This substantial gap arises because driving behavior requires long-chain reasoning (Perception→Prediction→Planning→Behavior), rather than simple pattern recognition. SFT, which mimics static, non-reasoning annotations, struggles to internalize the complex, multi-step logic necessary for high-level decision-making. In contrast, RL with Group Relative Policy Optimization (GRPO) enables the model to explicitly explore multiple reasoning paths, learning "how to think" rather than just "what to say." Furthermore, the complexity of autonomous driving, with its intricate logic and safety constraints, amplifies the SFT-RL gap. Unlike simpler VQA tasks, where this gap is minimal, the reasoning-intensive nature of behavior generation in driving tasks highlights the disproportionate benefits of RL’s exploration and self-correction capabilities.
Appendix L Case Study
@@
Figure 13: Perception task example comparing DriveLM and DriveRX. The question asks for the moving status of an object in the rear camera view (CAM_BACK). DriveLM hallucinates visual features, incorrectly stating that the object angles toward the side (highlighted in red). In contrast, DriveRX utilizes internal reasoning to correctly analyze the trajectory, noting that the object maintains a steady course without noticeable lateral shift (highlighted in blue), matching the Ground Truth “Going ahead”.
@@
Figure 15: Planning task example comparing Align-DS-V and DriveRX. The goal is to determine safe actions for the ego vehicle. Align-DS-V suggests a potentially risky action by advising to drift into the opposing lane to gain a better view past roadside barriers (highlighted in red). In contrast, DriveRX provides a safer and more prudent recommendation, suggesting to slightly offset within the lane to increase clearance and visibility while maintaining a safe following distance (highlighted in blue). This comparison highlights DriveRX’s ability to generate safer, lane-constrained planning decisions that align with the Ground Truth, avoiding the hazardous lane deviation proposed by Align-DS-V.
@@
Figure 17: Case study of DriveRX-Agent
Appendix M Experiment Prompt
M.1 Task-Specific Reward Prompt
@@
Appendix N Corruption Details